OpenStudy (anonymous):

find the derivative of 5sinx

geerky42 (geerky42):

Ok, what do you need to do?

OpenStudy (anonymous):

find the derivative of 5sinx

geerky42 (geerky42):

You know the derivative of sinx, right?

OpenStudy (anonymous):

ya its cosx

OpenStudy (anonymous):

i am confused since there is constant 5 in front?

geerky42 (geerky42):

Yeah, so \(\dfrac{\mathrm d}{\mathrm d x}cf(x) = c\dfrac{\mathrm d}{\mathrm d x}f(x) = cf'(x)\) So \(\dfrac{\mathrm d}{\mathrm d x}5\sin x = 5\dfrac{\mathrm d}{\mathrm d x}\sin x = \boxed{5\cos x}\)

geerky42 (geerky42):

Does that make sense?

OpenStudy (anonymous):

Thank you.

geerky42 (geerky42):

Ok, no problem